# for Large Language Models

### v1

@@ 3.3 Construction
InternLM2.5 7B 45.7 2.6 45.4 10.3 37.4 20.6 37.5 53.2 12.7 20.2 34.4 54.0
20B 51.9 4.5 51.2 14.4 42.9 24.0 - 26.6 14.9 19.7 34.9 53.9
Aya-Expanse 8B 51.2 9.5 33.8 7.8 50.8 26.2 - 41.1 21.5 26.8 45.6 51.6
32B 61.9 14.6 52.0 15.8 66.7 27.7 - 59.8 25.2 32.8 54.8 62.3
Gemma2 9B 63.0 9.1 53.9 16.6 72.0 23.9 - 61.4 27.2 33.2 57.5 61.9
27B 62.4 16.0 66.7 24.6 75.3 26.7 - 64.7 30.4 34.5 64.8 66.2
Llama3.1 8B 62.6 5.1 52.9 14.1 63.4 23.8 68.3 45.0 24.6 29.8 53.9 62.9
70B 76.2 11.1 69.7 29.8 79.7 35.8 57.4 44.3 31.1 35.1 64.5 68.2
Qwen2.5 7B 65.9 10.3 68.2 24.7 63.4 27.6 53.5 48.9 16.6 25.6 46.4 60.0
72B 80.8 34.1 78.6 45.5 77.8 39.4 80.6 61.8 25.8 33.3 60.4 66.9
DeepSeek-V3 671B 83.9 46.3 83.2 60.4 84.2 47.4 - 69.2 33.9 34.5 70.3 67.8
GPT-4o-mini - 79.1 23.6 78.7 37.0 76.9 34.1 82.1 70.9 30.3 33.9 67.7 67.6
@@ 4.1 Evaluation Setup
Evaluated Models We focus primarily on multilingual post-trained models and evaluate both open-source and proprietary language models [^4] Unless otherwise specified, all models discussed in this paper are post-trained versions., including Llama3.1 (Dubey et al., 2024), Qwen2.5 (Qwen Team, 2024), Gemma2 (Team et al., 2024), InternLM2.5 (Cai et al., 2024), Aya-Expanse (Dang et al., 2024), DeepSeek-V3 (DeepSeek-AI et al., 2024), and GPT-4o-mini (OpenAI, 2024). Models detailed descriptions are in Appendix C.
Inference Configuration We adopt greedy decoding on most tasks except the problem solving task, on which the temperature is set to $0.2$. The default chat template and system prompt of each model are applied. Detailed prompts are provided in Appendix D. Specifically, for instruction following tasks, the translated instructions are directly used as prompts. For reasoning tasks, we adopt the zero-shot native chain-of-thought (CoT) templates in LM-Evaluation-Harness (Gao et al., 2024). For other tasks, we use the original prompt templates provided in the corresponding repositories [^5] https://github.com/EleutherAI/lm-evaluation-harness
@@
Figure 4: Performance of several models on the science reasoning task and the domain translation task across different languages. Models show unbalanced multilingual capabilities on these tasks.
Figure 5: The proportion of tasks where larger models achieve a smaller GAP versus where smaller models perform better across different model families. larger models do not consistently have a smaller GAP, particularly in the Gemma2 family, where the smaller model outperforms the larger one in most cases.
@@ 4.2 Multilingual Benchmark Results
Table 6 shows the overall average performance of each model on each multilingual task. We take the top-performing ones and present their detailed performance on science reasoning and domain translation tasks, as illustrated in Figure 4. More detailed results are in Appendix E.
@@
As shown in Table 6, larger models consistently demonstrate enhanced multilingual capabilities across all domains, with few exceptions. However, the performance gap between English and non-English languages does not invariably diminish. We define GAP as the average performance gap between English and other languages: GAP=l≠enmax(s(en) - s(l), 0)n-1, where $s(l)$ denotes the score on the task with language $l$, and $n$ is the number of languages including English. As shown in Figure 5, when comparing models of different sizes, the proportion of larger models achieving smaller GAPs only slightly exceeds 0.5 for most model families except Qwen2.5. Gemma2-9B achieves smaller GAPs than Gemma2-27B in most tasks. These findings suggest that while scaling model size effectively improves overall multilingual performance, additional strategies may be needed to address the performance disparities across languages.
@@
Models’ reasoning capabilities vary significantly across languages (Figure 4). While models generally achieve better performance in dominant languages like English, their performance in other languages often lags behind. This disparity can be attributed to the fact that multilingual task execution depends not only on language-agnostic reasoning but also on language-specific capabilities such as comprehension and generation. Therefore, when operating in languages where the model has weaker proficiency, it becomes difficult to fully leverage its language-agnostic capabilities. Interestingly, we observed some unexpected patterns where certain models excel in specific non-dominant languages compared to English. For example, Qwen2.5 demonstrates superior performance in Korean over English on the science reasoning task. This counter-intuitive phenomenon warrants further investigation in future work.
Model performance exhibits systematic bias towards high-resource languages, with Gemma2 as a notable exception.
As shown in Figure 4, the performance curves of most models exhibit significant fluctuations across languages. High-resource languages such as French and Chinese consistently outperform low-resource languages like Telugu, Swahili, and Bengali. For instance, while DeepSeek-V3 achieves over 50% accuracy in science reasoning tasks for English and French, its performance drops notably to below 40% for Telugu. Our evaluation shows that almost all models have at least one language where they significantly underperform. This pattern can be partially attributed to development strategies - models like Aya-Expanse were not specifically optimized for the full range of languages in our evaluation. Notably, Gemma2 demonstrates unexpectedly balanced performance across most tasks (Figure 9), despite not being explicitly marketed as a multilingual model.
DeepSeek-V3 narrows the gap between open-source and closed-source models.
As demonstrated in Table 6 and Figure 9, GPT-4o-mini demonstrates strong multilingual capabilities across various tasks. While its performance matches that of Qwen2.5-72B on several tasks, it falls short of DeepSeek-V3, the leading open-source model in our evaluation. This finding suggests that state-of-the-art open-source models are becoming increasingly competitive with their closed-source counterparts. Due to budget constraints, our evaluation is limited to GPT-4o-mini among closed-source models, and a more comprehensive comparison would be valuable for validating this trend.
@@
We analyze the relationship between model’s English-to-X translation capability and other capabilities using Spearman correlation coefficients (the right panel of Figure 7). When calculating correlations between domain-specific translation and task performance, we exclusively use data from the corresponding domains. The analysis reveals that domain-specific translation performance generally exhibits stronger correlations with task performance compared to general translation capabilities. A notable exception is that in the rule-based instruction-following task, we observe an inverse scaling effect: larger LLMs produce lower-quality translations compared to their smaller counterparts. This occurs because larger LLMs are more likely to execute instructions rather than strictly perform translation, known as prompt injection.
Figure 6: F1 scores of correct answers in other languages to English across six tasks. Using English correctness as a reference, we treat other languages’ correctness as predictions to measure their agreement with English. Results show high agreement between English and other languages on these tasks.
Table 7: Different metrics are used to evaluate the En-X translation performance of selected models on specific domains.
@@
Figure 7: Left: Spearman Correlations between the performance of two models within the same family across different languages. Right: Spearman Correlations between the performance of models on general translation and domain translation.
We calculate Spearman correlation coefficients to analyze the performance similarity between models of the same family across different languages for each task. As shown in the left panel of Figure 7, models within the same family demonstrate strong correlations across various tasks, with most correlation coefficients exceeding 0.7. The exception is Llama3.1’s negative correlation in tool use tasks.
5 Analysis
5.1 High correctness agreement between English and other languages
Although sometimes similar performance can be achieved across different languages for certain tasks, the specific problems being addressed may vary significantly. To examine this language alignment, we compute F1 scores measuring the agreement between problem-solving correctness in English versus other languages. Using English correctness as the reference labels, Figure 6 shows the F1 scores of Llama3.1-70B and DeepSeek-V3 on six subtasks of BenchMAX. Both these strong multilingual models demonstrate high correctness agreement on most tasks, with average F1 scores exceeding 0.9. F1 scores for low-resource languages are notably lower than those for high-resource languages. This pattern is particularly pronounced in science reasoning tasks, suggesting these knowledge-intensive problems pose unique challenges for cross-lingual knowledge transfer.
5.2 Challenges in evaluating domain-specific translation quality
Domain-specific texts often contain substantial content that does not require translation, which leads to inflated spBLEU scores. To address this limitation, we explore alternative evaluation approaches: the edit-distance metric TER (Snover et al., 2006), the model-based metric xCOMET-XXL (Guerreiro et al., 2024), and a novel performance retention rate that compares downstream task performance between model self-translation and human translations. Table 7 presents these metric scores across selected tasks and languages. Traditional metrics prove unreliable for domain-specific translation evaluation. Both spBLEU and TER show extreme values in Science and Programming tasks due to large portions of unchanged text, failing to capture the quality of crucial translated segments. The model-based metric xCOMET shows high variance across different scenarios, particularly struggling with low-resource languages and specialized domains. Moreover, the performance retention rate doesn’t work as expected as many values are greater than 1, which fails to evaluate translations subtly. These findings highlight the need for specialized metrics for domain-specific translation evaluation, which we identify as an important direction for future research.
5.3 Machine-translated task data leads to inconsistent assessment of LLM capabilities
Evaluating models on machine-translated data can lead to both overestimation and underestimation of model performance. This inconsistency is evidenced by the retention rates exceeding 1.0 in Table 7, suggesting occasional overestimation. However, underestimation is more prevalent (Figure 8). Both Llama3.1-70B and Qwen2.5-72B achieve significantly higher scores on human-translated rule-based instruction following tasks across several languages, including French, Arabic, and Hungarian, with performance gaps exceeding 4 points in some cases.
Figure 8: The performance differences between using human-translated data and machine-translated data. The upper figure depicts the performance difference of Llama3.1-70B on the rule-based instruction-following task, while the lower figure presents the performance difference for Qwen2.5-72B on the same task.
5.4 Conventional understanding tasks are inadequate to evaluate multilingual capdbilities of LLMs
Prior work like Aya-Expanse (Dang et al., 2024) relies on conventional understanding tasks such as XCOPA and XWinograd for multilingual evaluation. On these metrics, Gemma2-9B achieves the best performance, followed by Aya-Expanse-8B, Llama3.1-8B, and Qwen2.5-7B. However, our evaluation through BenchMAX reveals a different pattern: Qwen2.5-7B demonstrates superior multilingual capabilities on generation tasks, while Aya-Expanse models show notably weaker performance on code generation tasks, as shown in Table 8. This discrepancy highlights the importance of comprehensive evaluation frameworks that incorporate both understanding and generation tasks to accurately assess the multilingual capabilities of LLMs.

### v2

@@ 3.3 Construction
InternLM2.5 7B 45.7 1.9 45.4 10.3 37.4 20.6 37.5 53.2 12.7 20.2 34.4 54.0
20B 51.9 3.3 51.2 14.4 42.9 24.0 - 26.6 14.9 19.7 34.9 53.9
Aya-Expanse 8B 51.2 6.4 33.8 7.8 50.8 26.2 - 41.1 21.5 26.8 45.6 51.6
32B 61.9 12.4 52.0 15.8 66.7 27.7 - 59.8 25.2 32.8 54.8 62.3
Gemma2 9B 63.0 9.8 53.9 16.6 72.0 23.9 - 61.4 27.2 33.2 57.5 61.9
27B 62.4 18.0 66.7 24.6 75.3 26.7 - 64.7 30.4 34.5 64.8 66.2
Llama3.1 8B 62.6 4.3 52.9 14.1 63.4 23.8 68.3 45.0 24.6 29.8 53.9 62.9
R1-Distill-Llama3.1 8B 49.7 3.5 62.8 23.8 46.9 28.1 - 37.2 12.2 20.8 13.5 23.1
Llama3.1 70B 76.2 13.2 69.7 29.8 79.7 35.8 57.4 44.3 31.1 35.1 64.5 68.2
Llama3.3 70B 85.2 17.0 74.0 34.7 83.8 42.6 50.4 42.5 31.5 33.6 63.5 65.0
R1-Distill-Llama3.3 70B 78.0 26.6 84.6 54.8 82.8 46.1 - 62.1 26.0 33.0 47.6 45.2
Qwen2.5 7B 65.9 8.5 68.2 24.7 63.4 27.6 53.5 48.9 16.6 25.6 46.4 60.0
R1-Distill-Qwen2.5 7B 46.7 3.0 69.3 37.3 56.1 28.4 - 27.7 6.8 16.3 17.0 27.3
Qwen2.5 32B 78.1 17.3 75.8 42.7 77.7 37.7 79.4 66.7 22.7 30.5 54.2 65.4
R1-Distill-Qwen2.5 32B 67.3 19.2 80.6 54.4 77.3 37.0 - 60.4 20.3 28.5 37.1 37.7
Qwen2.5 72B 80.8 36.9 78.6 45.5 77.8 39.4 80.6 61.8 25.8 33.3 60.4 66.9
DeepSeek-V3* 671B 83.9 59.8 83.2 60.4 84.2 47.4 85.2 69.2 33.9 34.5 70.3 67.8
GPT-4o-mini - 79.1 21.9 78.7 37.0 76.9 34.1 82.1 70.9 30.3 33.9 67.7 67.6
Table 6: Performance comparison across models on BenchMAX tasks, averaged over 17 languages. Bold numbers indicate the best performance in each column. "Func Compl." refers to Function Completion, "Prob. Solving" to Problem Solving, and "Multi Func." to Multiple Functions scenarios where models must select and call one function from multiple options. Models without results on the long context task do not support 128K context length. * DeepSeek-V3 is a 671B MoE model, with 37B activated for each token.
Step 2: Post-editing each sample by three distinct native-speaking annotators in almost all tasks.
To ensure high-quality dataset, we implement a multi-round annotation and verification process. 1) Each sample is given to three native-speaking annotators who are proficient in English and their native language. Considering the specialized nature of datasets like Science reasoning, annotators are required to hold at least a Bachelor’s degree. 2) Two automatic verifiers - rule-based verifiers and model-based verifiers - are used to assess the quality of human annotation. Rule-based verifiers ensure the satisfaction of constraints for certain tasks, such as the rule-based instruction following task. For model-based verifiers, we utilize the GEMBA-SQM prompt Kocmi and Federmann (2023) and employ Qwen2.5-72B-Instruct, a powerful multilingual model, to estimate the quality of translations. Along with providing an overall score, the model offers detailed explanations of translation errors as feedback to annotators. Samples that do not pass the rule-based verifier or score below a predefined threshold are identified as failed, and refined in subsequent iterations. Each manually annotated dataset undergoes at least three iterations.
Step 3: Selecting the final translation version by LLMs.
Initially, we ask a fourth annotator, uninvolved in the annotation process, to choose the final version from the results revised by three individuals. Intriguingly, the selection by the fourth annotator exhibited a strong position bias, often favoring the initial annotation. This preference could be attributed to the uniformly high quality of translations, resulting in minimal discernible differences among them.
Debiasing for human annotators is costly in terms of both time and finance, because three translations encompass all permutations of six. Consequently, we employ GPT-4o-mini to select the final translation, as it is a powerful and balanced LLM across different languages. In particular, following Li et al. (2024), we adapt the LLM-Judge system instruction (see Appendix D) to suit pairwise translation evaluation. We shuffle the positions of three translations and conduct two battles to select a final version. In each battle between two translations, we perform two judgments by swapping their positions and determine one winner. The winner of the first two translations then battles against the third translation, determining the final winner.
@@ 4.1 Evaluation Setup
Evaluated Models We focus primarily on multilingual post-trained models and evaluate both open-source and proprietary language models [^4] Unless otherwise specified, all models discussed in this paper are post-trained versions., including Llama3.1 Dubey et al. (2024), Qwen2.5 Qwen Team (2024), Gemma2 Team et al. (2024), InternLM2.5 Cai et al. (2024), Aya-Expanse Dang et al. (2024), DeepSeek-R1-Distill-Llama Guo et al. (2025), DeepSeek-R1-Distill-Qwen, DeepSeek-V3 DeepSeek-AI et al. (2024), and GPT-4o-mini OpenAI (2024). Models’ detailed descriptions are in Appendix C.
Inference Configuration We adopt greedy decoding for most tasks, except for the problem solving task, where the sampling temperature is set to $0.2$. The default chat template and system prompt of each model are applied. Detailed prompts are provided in Appendix D. For reasoning tasks, we adopt the zero-shot native chain-of-thought templates in LM-Evaluation-Harness Gao et al. (2024). For other tasks, we use the prompt templates provided in corresponding repositories [^5] https://github.com/EleutherAI/lm-evaluation-harness
@@
Figure 4: Taking two tasks as examples, models exhibit unbalanced multilingual capabilities. We show performance of several models on the science reasoning task and the domain translation task across different languages.
Figure 5: Larger models do not consistently have a smaller GAP. Each row shows proportions of tasks where the larger model achieves a smaller GAP versus where the smaller model performs better.
@@ 4.2 Multilingual Benchmark Results
Table 6 shows the overall average performance of each model on each multilingual task. More detailed results are in Appendix E.
@@
As shown in Table 6, larger models consistently demonstrate enhanced multilingual capabilities across all domains, with few exceptions. However, the performance gap between English and non-English languages does not invariably diminish. We define GAP as the average performance gap between English and other languages: GAP=∑l≠enmax(s(en) - s(l), 0)n-1, where $s(l)$ denotes the score on the task with language $l$, and $n$ is the number of languages including English. As shown in Figure 5, when comparing models of different sizes, the proportion of larger models achieving smaller GAPs only slightly exceeds 0.5 for most model families. Gemma2-9B achieves smaller GAPs than Gemma2-27B on most tasks. These findings suggest that while scaling model size effectively improves overall multilingual performance, additional strategies may be needed to address the performance disparities across languages.
@@
As illustrated in the left plot of Figure 4, models’ reasoning capabilities vary significantly across languages. Generally, models achieve better performance in dominant languages like English than in other low-resource languages. This disparity can be attributed to the fact that multilingual task execution depends not only on language-agnostic reasoning but also on language-specific capabilities such as comprehension and generation. Therefore, when operating in weak languages, it becomes difficult for a model to fully leverage its language-agnostic capabilities. Interestingly, we observe an unexpected pattern where certain models excel in specific non-dominant languages compared to English on some tasks. For example, Qwen2.5 demonstrates superior performance in Korean over English on the science reasoning task. We hypothesize that Qwen2.5’s training data includes a relatively high proportion of Korean content in scientific or reasoning-related domains. This counter-intuitive phenomenon warrants further investigation in future work.
Figure 6: Left: The translation performance is positively correlated with other multilingual performance. Spearman Correlations are computed between the performance on general/domain translation and the specific task. Right: Models in the same family have similar language performance pattern. We compute the Spearman Correlations between the performance of two models (Llama3.1 8B vs 70B, Qwen2.5 7B vs 72B) across different languages.
Model performance exhibits systematic bias towards high-resource languages.
As shown in Figure 4, the performance curves of most models exhibit significant fluctuations across languages. High-resource languages such as French and Chinese consistently outperform low-resource languages like Telugu, Swahili, and Bengali. For instance, while DeepSeek-V3 achieves over 50% accuracy in science reasoning tasks for English and French, its performance drops notably to below 40% for Telugu. This pattern can be partially attributed to development strategies - models like Aya-Expanse were not specifically optimized for the full range of languages in our evaluation. Unexpectedly, Gemma2 exhibits relatively balanced performance across most tasks (Figure 9), despite not being explicitly marketed as a multilingual model.
@@
We analyze the relationship between model’s English-to-X translation capability and other capabilities using Spearman correlation coefficients (the left panel of Figure 6). When calculating correlations between domain-specific translation performance and task performance, we exclusively use data from the corresponding domains. The analysis reveals that domain-specific translation performance generally exhibits stronger correlations with task performance compared to general translation capabilities. A notable exception is that in the rule-based instruction-following task, we observe an inverse scaling effect: larger LLMs produce lower-quality translations compared to their smaller counterparts. We find that larger LLMs are more likely to execute instructions rather than strictly perform translation, known as prompt injection.
We calculate Spearman correlation coefficients to analyze the performance similarity between models of the same family (excluding R1-distilled models) across different languages for each task. As shown in the right panel of Figure 6, models within the same family show strong correlations across various tasks, with most correlation coefficients exceeding 0.7.
R1-distilled models exhibit enhanced multilingual reasoning and code generation capabilities, but some other capabilities, especially translation, are noticeably degraded.
As illustrated in Table 6, the performance of R1-Distill-Llama3.3-70B is comparable to DeepSeek-V3 in reasoning and code generation tasks, and is stronger than Llama-3.3-70B-Instruct. However, other capabilities like instruction following of 7B/8B models exhibit degradation to some extent. They tend to generate repeated tokens in the reasoning process when using non-English languages. The translation capabilities of both large and small distilled models decline dramatically. In addition to repeated generation, we also observe a frequent phenomenon of code-switching in translations.
5 Analysis
Llama3.1-70B Qwen2.5-72B
Translated by GT 4o-mini Ours GT 4o-mini Ours
Rule-based 66.9 53.5 76.2 71.5 57.2 80.8
Func Compl. 47.8 68.2 69.7 50.4 75.5 78.6
Science 33.7 35.1 35.8 36.9 37.8 39.4
Multi Func. 23.0 43.7 44.3 26.7 61.3 61.8
Table 7: Our pipeline provides a more accurate assessment of the multilingual performance, compared to naive translations by Google Translate(GT) and GPT-4o-mini(4o-mini), respectively.
5.1 Our pipeline provides a more accurate assessment of models’ performance
We naively translate a subset of tasks from English to other 16 languages by Google Translate and GPT-4o-mini, and then evaluate two models using this task data. We directly translate sources by Google Translate as it doesn’t support constraints, and use appropriate prompts with constraints to request GPT-4o-mini. The results in Table 7 show that models achieve higher scores generally on our translated tasks compared to naive machine-translated ones. Google Translate lacks the flexibility to handle diverse constraints and specific domains, while GPT-4o-mini does not always perform translation task, especially on instruction data. This indicates that naive machine translation underestimates LLMs’ capabilities, whereas our data provides a more accurate assessment.
5.2 BenchMAX is adequate to evaluate multilingual capabilities of LLMs
Prior work like Aya-Expanse Dang et al. (2024) relies on conventional understanding tasks such as XCOPA and XWinograd for multilingual evaluation. On these metrics, Gemma2-9B achieves the best performance, followed by Aya-Expanse-8B, Llama3.1-8B, and Qwen2.5-7B. However, our evaluation through BenchMAX reveals a different pattern: Qwen2.5-7B demonstrates superior multilingual capabilities on generation tasks, while Aya-Expanse models show notably weaker performance on code generation tasks, as shown in Table 8. This discrepancy highlights the importance of comprehensive evaluation frameworks that incorporate both understanding and generation tasks to accurately assess multilingual capabilities of LLMs.
Model Rule-based Func Compl. Discriminative
Qwen2.5-7B No.1 No.1 No.4
Llama3.1-8B No.3 No.3 No.3
Aya-Expanse-8B No.4 No.4 No.2
Gemma2-9B No.2 No.2 No.1
Table 8: Rankings of the models in generative tasks in BenchMAX differ from that in discriminative tasks, indicating the importance of both types of tasks. The ranking in discriminative tasks is from Dang et al. (2024).
5.3 BenchMAX reveals the challenges in domain-specific translation evaluation
Domain-specific texts often contain substantial content, such as code, that does not require translation, leading to inflated spBLEU scores. To address this limitation, we explore alternative evaluation approaches: the edit-distance metric TER Snover et al. (2006), the model-based metric xCOMET-XXL Guerreiro et al. (2024), and a novel performance retention rate that compares downstream task performance between model self-translation and human translations. Table 9 presents these metric scores across selected tasks and languages. Traditional metrics prove unreliable for domain-specific translation evaluation. Both spBLEU and TER show extreme values in Science and Programming tasks due to large portions of unchanged text, failing to capture the quality of crucial translated segments. The model-based metric xCOMET scores show inconsistency, from 18 to 96, across different scenarios, particularly struggling with low-resource languages and specialized domains. Moreover, the performance retention rate doesn’t work as expected as many values are very close, which fails to evaluate translations directly and subtly. These findings highlight the need for specialized metrics for domain-specific translation evaluation, which we identify as an important direction for future research.
Figure 7: Advanced models show high consistency between English and other languages across six tasks.
5.4 High consistency between the questions answered correctly/incorrectly in English and in other languages
Although sometimes similar performance can be achieved across different languages for certain tasks, the specific problems being addressed may vary significantly. To examine the language alignment, we compute the consistency between the problem-solving correctness in English versus other languages. Consistency is calculated as the proportion of predictions where a model’s output is correct or incorrect in both languages, out of all evaluated samples. Figure 7 presents the consistency between English and languages, based on results of Llama3.1-70B and DeepSeek-V3 on six subtasks of BenchMAX. Both these strong multilingual models demonstrate high consistency on most tasks, with most scores exceeding 0.75. Agreement for low-resource languages are notably lower than those for high-resource languages. Low consistency is also pronounced in science reasoning tasks, suggesting these knowledge-intensive problems pose unique challenges for cross-lingual knowledge transfer.
@@ 5.3 BenchMAX reveals the challenges in domain-specific translation evaluation
Table 9: There exists challenges in domain-specific translation evaluation. The table presents different metric scores of the En-X translation of selected models on specific domains.
5.5 Self-bias is inevitable in the model-based instruction following tasks
Applying model-based evaluation exhibits self-bias, where the judge model prefers the outputs of itself or models from the same family Li et al. (2024); Xu et al. (2024). We further adopt GPT-4o-mini as the judge model, and compute the win-rate against the baseline model GPT-4o. Figure 8 shows that DeepSeek-V3 strongly favors its own outputs, while GPT-4o-mini prefers GPT-4o’s outputs. Nevertheless, the win-rates of other models judged by the different judges are comparable, and the rankings are fairly consistent.
Figure 8: Self-bias is inevitable in model-based instruction following evaluation. DeepSeek-V3 prefers its own outputs, while GPT-4o-mini tends to prefer GPT-4o’s outputs. The win-rates of evaluated models are judged by DeepSeek-V3 and GPT-4o-mini.
5.6 Comaparing open-source and closed-source models on BenchMAX
As demonstrated in Table 10, although GPT-4o-mini and GPT-4o demonstrates strong multilingual capabilities across various tasks, they fall short of DeepSeek-V3, the leading open-source model in our evaluation. This suggests that state-of-the-art open-source models are becoming competitive with their closed-source counterparts. Due to budget constraints, our evaluation of closed-source models is limited to GPT-4o-mini and GPT-4o on some of tasks. A more comprehensive comparison would be valuable for further validating this trend.
Model Rule-based Func Compl. Science
GPT-4o-mini 79.1 78.7 34.1
GPT-4o 80.8 80.1 45.6
DeepSeek-V3 83.9 83.2 47.4
Table 10: The leading open-source model, DeepSeek-V3, bridges the gap to the closed-source models. We compare DeepSeek-V3 and GPT-4o on some of tasks.